If you have 500 mL of a 0.10 M solution of the acid, what mass of the corresponding sodium salt of the conjugate base do you need to make the buffer with a pH of 2.06 (assuming no change in volume)?

chlorous acid (HClO2HClO2), pKa=1.95

I get 2.89g. Not correct.
Answer #1

no of moles of HClO2 = molarity* volume in L

                                 = 0.1*0.5 = 0.05moles

PH   = PKa + log[NaClO2]/[HClO2]

2.06 = 1.95 + log[NaClO2]/[HClO2]

log[NaClO2]/[HClO2] = 2.06-1.96

log[NaClO2]/[HClO2]    = 0.1

[NaClO2]/[HClO2]     = 1.26

[NaClO2]   = 1.26*[HClO2]

                = 1.26*0.05

                 = 0.063moles

mass of NaClO2   = no of moles* gram molar mass

                           = 0.063*90.5 = 5.7g >>>>answer